20% of students in a class go to professor during office hours. of those who go 30% seek minor clarification. 70% seek major clarification. a) what is the probability a student goes for minor clarification? b) what is the probability a student goes for major clarification? 


1
Expert's answer
2021-09-21T02:39:01-0400

Let A be the event that students go to professor and B the event that minor clarifications are sought. BC is the event that major clarifications are sought.

The information in this question can be summarized as follows,

P(A)=0.2

P(B|A)=0.3

P(BC|A)=0.7

To determine P(B), the property P(BC AND A)= P(BC)-P(B AND A) is applied, then obtain P(B)=1-P(BC).

P(A AND B)=P(B AND A)=P(B|A)*P(A)

=0.3*0.2

=0.06

P(BC AND A)=P(A AND BC)=P(BC|A)*P(A)

=0.7*0.2

=0.14

Therefore, we have

0.14=P(BC)-0.06

P(BC)=0.2, hence P(B)=1-0.2=0.8

a.

We find P(A|B)=P(A AND B)/P(B)

=0.06/0.8

=6/80

=3/40

b.

We find P(A|BC)=P(A AND BC)/P(BC)

=0.14/0.2

=7/10